OpenStudy (anonymous):

Write an equation of a line that passes through the point (8, 4) and is parallel to the line y = 4x + 2. a. y = 4x - 28 b. y = 4x + 28 c. y = 1/4 x - 2 d. y = 1/4 x + 2

OpenStudy (anonymous):

@ranga ?

OpenStudy (ranga):

Answer these questions in order and you would have solved the problem. 1) What is the slope of y = 4x + 2 ? 2) Parallel lines have the same slope. So what is the slope of a line parallel to y = 4x + 2 ? 3) The equation of a line in slope intercept form is y = mx + b Substitute the slope m from (2) above. The parallel line passes through the point (8,4) Substitute x = 8 and y = 4 and solve for b.

OpenStudy (anonymous):

THe slope is going to be the same, the only thing that will change is y intercept. Say, y = 4x + b and plug (8, 4) into this. y = 4x + b ---> 4=4(8)+b ---->4=32+b ---> b=-28 so y = 4x - 28. It is A.

OpenStudy (anonymous):

thanks! cx